A stock has an expected return of 11 percent, its beta is 1.60, and the expected return on the market is 9 percent. What must the risk-free rate be? (Round your answer to 2 decimal places. (e.g., 32.16))

  

  Risk-free rate %
Answer #1

let risk free rate = r

11 = r+1.6*(9-r)

11-1.6*9 = -0.6r

r= -(11-1.6*9)/0.6 = 5.67 %

riskfree rate = 5.67 %